OpenStudy (anonymous):

Which expressions show the expanded form for 907.03? Choose all answers that are correct. A. 9 × 100 + 7 × 1 + 0 × 1/1 + 0 × 1/10 B. 9 × 100 + 0 × 10 + 7 × 1 + 0 ×1/10+ 3 × 1/100 C. 9 × 100 + 7 × 10 + 3 × 1/100 D. 9 × 100 + 7 × 1 + 3 × 1/100

Directrix (directrix):

Go through the options and check them. For example, option A: 9 × 100 + 7 × 1 + 0 × 1/1 + 0 × 1/10 = ( 9 × 100) + ( 7 × 1) + ( 0 × 1/1) +( 0 × 1/10) = 900 + 7 + 0 + 0 = 907 Do that for option B and see what you get.

OpenStudy (anonymous):

so the answer is A. and B.?